OpenStudy (anonymous):

The inverse of the function y = 7x + 5 is y=x+5/7 ???

OpenStudy (anonymous):

is it true or false

OpenStudy (tyteen4a03):

x = 7y + 5 7y = x - 5 y = (x-5)/7 Swap the variables and work out the equation so that y becomes the subject.

OpenStudy (anonymous):

so true??? right

OpenStudy (anonymous):

nope.

OpenStudy (tyteen4a03):

@johana My solution was y = (x-5)/7, not y = (x+5)/7. You had the signs wrong.

OpenStudy (anonymous):

@tyteen4a03 says, swap the variables x>y as he did and rearrange to solve for y.

OpenStudy (anonymous):

so to inverse .. this sing have to change ??? so the answer is false ok thank u :))

OpenStudy (tyteen4a03):

No, to inverse you need to swap the variables and change the equation so it becomes the form y = something again.

OpenStudy (anonymous):

look it a true or false question i dont have to do any work just to know if it right or wrong that it

OpenStudy (anonymous):

yes, the answer is false but @tyteen4a03 was just trying to help you understand it better.

OpenStudy (anonymous):

im not trying to be mean or anything

OpenStudy (anonymous):

welll thank u tyteen4a03

OpenStudy (anonymous):

and jayds

OpenStudy (tyteen4a03):

@johana If I just told you the answers, you will not be able to do future questions like this one. Teaching a man how to fish is better than just giving him a fish.

OpenStudy (anonymous):

@ tyteen4a03 thank u for teaching it to me and sorry if my typing sound or in any way rude to u ok im just not myself today it been a long day and im at the edge off just quitting on flvs because math is jut not for me

OpenStudy (anonymous):

@ tyteen4a03 but i wanna graduate and prof my parent wrong my parent wrong u know

OpenStudy (tyteen4a03):

@johana This happens to me a lot of times too, just go grab some snacks and continue later ;P